OpenStudy (anonymous):

Simplify the expression. 6k^-6b^6

OpenStudy (anonymous):

\[6k ^{-6b}^{6}\] right?

OpenStudy (anonymous):

i dont know how to read that lol hold on

OpenStudy (anonymous):

ok use the rule \[x ^{-a}=\frac{ 1 }{ x ^{a} }\]

OpenStudy (anonymous):

so its 6k-6=1/b^6?

OpenStudy (anonymous):

the negative sign over k or b?

OpenStudy (anonymous):

over k

OpenStudy (anonymous):

nnvm i got it xD

OpenStudy (anonymous):

ok apply the rule for k not for b